The width and length of a rectangle are consecutive integers. If the perimeter of the rectangle is 166 inches, find the width and length of the rectangle.
Answer by sudhanshu_kmr(1152)


2( x + x+1 ) = 166
=> 4x + 2 = 166
=> 4x = 166-2 = 164
=> x = 41
so, dimension is: 41 x 42